OpenStudy (anonymous):

A hairdresser is deciding where to open her own studio. if the hairdresser chooses location A, she will pay $1200 per month in rent and will charge $45 per haircut. if she chooses location B, she will pay $1800 per month in rent and will charge $6o per haircut. how many haircuts would she have to give in one month to make the same profit at either location?

OpenStudy (anonymous):

y=45x-1200 y=60x-1800 45x-1200=60x-1800 15x=600 x=40

OpenStudy (anonymous):

How do you solve though??